a clock with metal pendulum beating seconds keep correct time at zeroDegree calsius. If it loses 12.5s a day at 25degreecalcius,the coefficient of linear expansion of metal pendulum is 

Grade:12th pass

1 Answers

Arun
5 years ago
Now, we have

Time lost in a day = time lost per second x number of seconds in a day
so,

dt = (1/2).α.T x 86400
or
12.5 = (1/2).α.25 x 86400
so,
α = (12.5 x 2) / (25 x 86400) 
thus, 
coefficient of linear expnsion is
α = (1/86400) oC-1